OpenStudy (anonymous):

A transmission line carries a current of 100 A. in east to west direction. Find magnitude of the magnetic field due to current at a distance 1m below the wire: (a) 2 × 10–5 T (b) 2 × 10+5 T (c) 105 T (d) 10–5 T

OpenStudy (anonymous):

B=\[\mu\]I/(2piR)(For an infinitely long wire) =(4pi/10^7*100)/(2pi*1) (\[\mu\]=4pi/10^7) =2*10^-5T